OpenStudy (anonymous):

A model rocket is launched from a roof into a large field. The path of the rocket can be modleled by the equation y = -0.8x^2 + 12x + 25.8 where x is the horizontal distance, in meters, from starting point on the roof and y is the height, in meters, of the rocket above the ground. How far horizontally from its starting point will the rocket land? Round your answer to the nearest hundredth?

OpenStudy (anonymous):

please don't type a reply and leave like other people I actually need help kind of walking through it so I can understand

OpenStudy (ranga):

When the rocket lands on the ground, its height above the ground, y, will be zero. Set y = 0 and solve for x.

OpenStudy (jdoe0001):

there's an assumption you've covered solving quadratic equations, since this is a quadratic

OpenStudy (anonymous):

Yes kind of but I'm am not sure or have a clue how to do this one that's why I closed the last question because it was just an equation you gave me when I need help completing the steps

OpenStudy (jdoe0001):

hmmm is really all you need to solve it, just set y = 0,and solve for "x" judging from the coefficients, I'd think you may need to use the quadratic formula

OpenStudy (anonymous):

I am not exactly sure what the quadratic formula is, I did all the homework for this unit and I am just dead in the water on this question for some reason

OpenStudy (jdoe0001):

meaning you'd need to cover the section on quadratics first

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=1Pva-Iv43Nc <--- solving by factoring and using the quadratic formula

OpenStudy (ranga):

\[\Large x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\\a = -0.8, b = 12, c = 25.8\]

OpenStudy (anonymous):

so 25.80 would be my answer if I worked it out right?

OpenStudy (anonymous):

ranga how exactly do I give you a medal?

OpenStudy (ranga):

How did you get 25.8 without plugging a,b and c values into the quadratic formula given above?

OpenStudy (anonymous):

wait a sec ill work it out

OpenStudy (ranga):

\[x = \Large \frac{ -12 \pm \sqrt{12 ^{2} - 4 * (-0.8) * 25.8} }{ 2* (-0.8) }\]

OpenStudy (anonymous):

-12-15.051/.16?

OpenStudy (anonymous):

My choices are or 25.80, or 37, or 17.24, or 16.91

OpenStudy (ranga):

\[x = \frac{ -12 \pm 15.052 }{ -1.6 }.~~~x = \frac{ -12 + 15.052 }{ -1.6 }~~or~~x= \frac{ -12 - 15.052 }{ -1.6 }\]

OpenStudy (ranga):

The first solution will give negative x which is not a valid solution. Try the second solution: \[x = \frac{-12-15.052}{-1.6} = \frac{-27.052}{-1.6} = 16.91\]